OpenStudy (anonymous):

__ Where P is (-2, 6), R is (-10, 4), and segment RT is a diameter of P, find the coordinates of T.

OpenStudy (anonymous):

*___ RT

myininaya (myininaya):

are you sure this is the question if p is a point then how can RT be a diameter of p this question makes no sense i'm sorry

OpenStudy (saifoo.khan):

Cool, Myinina is here, so no tension!

OpenStudy (anonymous):

p is a circle

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

hehe

OpenStudy (anonymous):

im going to post the question

OpenStudy (anonymous):

can you figure it out?

myininaya (myininaya):

okay so P is the center

OpenStudy (anonymous):

yes

myininaya (myininaya):

so we need to find the distance from R to P and that will be the distance from P to T so we have the length of RP is \[\sqrt{(-10-(-2))^2+(4-6)^2}=\sqrt{64+4}=\sqrt{68}=2\sqrt{17}\] so the length of PT is the same so the length of the dimeter is PT+RP or you could say 2*RP 2*RP is \[2(2\sqrt{17})=4\sqrt{17}\]

myininaya (myininaya):

ok we aren't done yet k

OpenStudy (anonymous):

ok...

myininaya (myininaya):

i'm going to post an attachment let me finish it first k?

OpenStudy (anonymous):

alright

myininaya (myininaya):

so x=6

myininaya (myininaya):

y=8

myininaya (myininaya):

i will post what i did k?

OpenStudy (anonymous):

cool!

myininaya (myininaya):

(6,8)

myininaya (myininaya):

i made two equations from the info that was given

OpenStudy (anonymous):

thanx! you really helped me! thats awesomeee!